# Diffusing DeBias: Synthetic Bias Amplification for Model Debiasing

### v3

@@ Abstract
Deep learning model effectiveness in classification tasks is often challenged by the quality and quantity of training data whenever they are affected by strong spurious correlations between specific attributes and target labels. This results in a form of bias affecting training data, which typically leads to unrecoverable weak generalization in prediction. This paper aims at facing this problem by leveraging bias amplification with generated synthetic data: we introduce Diffusing DeBias (DDB), a novel approach acting as a plug-in for common methods of unsupervised model debiasing exploiting the inherent bias-learning tendency of diffusion models in data generation. Specifically, our approach adopts conditional diffusion models to generate synthetic bias-aligned images, which replace the original training set for learning an effective bias amplifier model that we subsequently incorporate into an end-to-end and a two-step unsupervised debiasing approach. By tackling the fundamental issue of bias-conflicting training samples memorization in learning auxiliary models, typical of this type of techniques, our proposed method beats current state-of-the-art in multiple benchmark datasets, demonstrating its potential as a versatile and effective tool for tackling bias in deep learning models.
footnotetext: Correspondence: vito.paolo.pastore@unige.it
Figure 1: Comparison of bias amplifier training approaches. (a) Traditional methods that use real biased datasets lead to the memorization of biased-conflicting samples, resulting in suboptimal auxiliary models. (b) Our proposed approach leverages conditional diffusion models to learn class-specific biases and amplify them into synthetic images. Such generations can replace the original training set for learning an effective Bias Amplifier, eliminating memorization effects by avoiding exposure to real data.
@@ 1 Introduction
Deep learning models have shown impressive results in image classification, yet their success depends heavily on training data quality and representativeness. When a deep neural network is trained on a biased dataset, so-called “shortcuts”, corresponding to spurious correlations (i.e., biases) between irrelevant attributes and labels, are learned instead of semantic class attributes, affecting model generalization and impacting test performances. In other words, the model becomes biased towards specific training sub-populations presenting biases, known as bias-aligned, while samples not affected by the bias are referred to as bias-conflicting or unbiased.
In this scenario, several debiasing methods have been proposed [1, 2, 3, 4, 5], addressing the problem under different perspectives. Notably, the majority of available training samples are bias-aligned in a biased dataset (typically, $95\%$ or higher [6]). As such, common supervised debiasing strategies involve reweighting the training set to give more importance to the few available bias-conflicting data [5], for instance, by upsampling [7] or up-weighting [8] them. Conversely, unsupervised debiasing methods, which assume no prior availability of bias information, typically involve an auxiliary model to estimate and convey information on bias for reweighting the training set, employing their signal to guide the training of a debiased target model, while diminishing its tendency to learn bias shortcuts.
@@
To this end, there are several existing works [6, 1, 5, 9] exploiting intentionally biased auxiliary models, trained to overfit bias-aligned samples, thus being very confident in both correctly predicting bias-aligned training samples as well as misclassifying bias-conflicting ones [5]. Such a distinct behavior between the two subpopulations is required to convey effective bias supervisory signals for target model debiasing, including gradients or per-sample loss values, which, e.g., should ideally be as high as possible for bias-conflicting and as low as possible for bias-aligned [6]. However, such desired behavior is severely hindered by bias-conflicting training samples, despite their low number. In fact, the low cardinality of this population impedes a suitable modeling of the correct data distribution, yet, they are sufficient to harm the estimation of a reliable bias-aligned data distribution, despite the latter having much higher cardinality. First, bias-conflicting samples in the training set act as noisy labels, affecting the bias learning for the auxiliary models [5]. Furthermore, these models tend to overfit and memorize bias-conflicting training samples within very few epochs, failing to provide different learning signals for the two subpopulations [10]. Despite available strategies, including the usage of the Generalized Cross Entropy (GCE) loss function [6, 3], ensembles of auxiliary models [1, 5], or even ad hoc solutions tailored to the specific dataset considered (e.g., training for only one epoch as in [11]), how to define a protocol for effective and robust auxiliary model training for unsupervised debiasing approaches remains an open issue.
@@
Remarkably, recent works [3, 5, 1] underline how the target model generalization is profoundly impacted by the protocol used to train such an auxiliary model. In principle, if one could remove all bias-conflicting samples from the training set, it would be possible to train a bias-capturing model robust to bias-conflicting samples’ memorization and interference, and capable of conveying ideal learning signals [5].
@@
To this aim, our intuition is to elude these drawbacks by skipping the use of the actual training data in favor of another set of samples synthetically generated by an adequately trained diffusion model. Specifically, we propose to circumvent bias-conflicting interference on auxiliary models leveraging fully biased synthetic data exploiting intrinsic properties of Conditional Diffusion Probabilistic Models (CDPMs). We aim to learn a per-class biased image distribution, from which we sample an amplified synthetic bias-aligned subpopulation. The resulting synthetic bias-aligned samples are then exploited for training a Bias Amplifier model instead of the original actual training set, solving the issue of bias-conflicting overfitting and interference by construction, as the bias-capturing model does not see the original training set at all. We refer to the resulting approach as Diffusing DeBias (DDB). DDB’s intuition is indirectly supported by recent works studying the problem of biased datasets for image generation tasks  [12, 13], and showing how diffusion models’ generations can be biased, as a consequence of bias present in the training sets (See Sec. 2).
@@
This tendency can be seen as an undesired behavior of generative models, impacting the fairness of generated images, and recent papers have started to propose solutions for decreasing diffusion models’ inclination to learn biases [14, 13, 15]. However, we claim that what is considered a problem for generative tasks can be turned into a feature for image classification model debiasing, as it allows the training of a robust auxiliary model that can ideally be plugged into various debiasing schemes, solving bias-conflicting training samples memorization, by construction (see Figure 1). For this reason, we refer to DDB as a plug-in for debiasing methods in image classification. To prove its effectiveness and versatility, we incorporate DDB’s bias amplifier into two different debiasing frameworks, named Recipe I and Recipe II, able to: (i) extract subpopulations further used as pseudo-labels within the popular G-DRO [8] algorithm (Recipe I, Sec. 3.3.1), and (ii) provide a loss function for the training set to be used within an end-to-end method (Recipe II, Sec. 3.3.2). Both approaches outperform the state-of-the-art on popular benchmark datasets. To summarize, our main contributions can be summarized as follows:
@@
  • We introduce DDB, a novel and effective unsupervised debiasing framework that exploits intrinsic properties of CDPMs to learn per-class bias-aligned distributions in biased datasets, with unknown bias information. The resulting CDPM is used for generating synthetic bias-aligned images later employed for training a robust Bias Amplifier model, hence avoiding bias-conflicting training sample memorization and interference (Sec. 3).
@@
  • We propose the usage of DDB as a versatile plug-in for debiasing approaches, designing two recipes: one based on a two-step procedure (Sec. 3.3.1), and the other as an end-to-end (Sec. 3.3.2) algorithm.
  • Both proposed approaches prove to be effective, surpassing state-of-the-art results by a significant margin (Sec. 4) across four different benchmark datasets.
@@ 2 Related Work
In this section, we provide a broad categorization of model debasing works.
@@
Supervised Debiasing. Supervised methods exploit available prior knowledge about the bias (e.g., bias attribute annotations), indicating whether a sample exhibits a particular bias (or not). A common strategy consists of training a bias classifier to predict such attributes so that it can guide the target model to extract unbiased representations [16, 17]. Alternatively, bias labels can be thought of as pre-defined encoded subgroups among target classes, allowing for the exploitation of robust training techniques such as G-DRO [8]. At the same time, bias information can be employed to apply regularization schemes aiming at forcing invariance over bias features [18, 19].
@@
Unsupervised Debiasing. It considers bias information unavailable (the most likely scenario in real-world applications), thus having the highest potential impact among the described paradigms. Among such methods, a popular strategy consists in involving auxiliary models at any level of the debiasing process. We categorize these methods into two-step  [11, 9, 4]), and end-to-end  [6, 20]), in the following paragraphs.
@@
End-to-end approaches. Here, bias mitigation is performed online with dedicated optimization objectives. Employing techniques such as ensembling and noise-robust losses have been proposed, though they require careful tuning of all the hyperparameters, which may require the availability of validation sets with bias annotations to work. In Learning from Failure (LfF), Nam et al. [6] assume that bias features are learned earlier than task-related ones: they employ a bias-capturing model to focus on easier (bias-aligned) samples through the GCE loss [21]. At the same time, a debiased network is jointly trained to assign larger weights to samples that the bias-capturing model struggles to discriminate. In DebiAN, Li et al. [20] jointly train an auxiliary model (the discoverer) and a classifier (the target model) in an alternate scheme so that the discoverer can identify bias learned by the classifier during training, optimizing an ad-hoc loss function for bias mitigation.
@@
Two-step methods. Here, the auxiliary model is employed in a first step to identify bias in training data, generally through pseudo-labeling, while the second stage consists in actual bias mitigation exploiting the inferred pseudo-labels. In [11], the auxiliary model is trained with standard ERM and then used in inference on the training set, considering misclassified samples as bias-conflicting and vice-versa: debiasing is brought on by up-sampling the predicted bias-conflicting samples. Notably, they must rely on a validation set with bias annotations to stop the auxiliary model’s training after a few epochs to avoid memorization and to tune the most effective upsampling factor. In [1], a set of auxiliary classifiers is ensembled (bootstrap ensembling) into a bias-commmittee, proposing that debiasing can be performed using a weighted ERM, where the weights are proportional to the number of models in the ensemble misclassifying a certain sample. Finally, a recent trend involves the usage of vision-language models to identify bias, further exploiting bias predictions to mitigate biases in classification models [22, 23, 24].
@@
Among the described methods, a subset exploit the auxiliary model for bias amplification ([5, 6, 9]) under the general assumption that a bias amplification model would either misclassify or be less confident on bias-conflicting samples, with final debiasing performance strongly dependent on such behavior, and specifically, on how strongly the auxiliary model captures training bias [3, 5]. However, this fundamental assumption only holds if bias-conflicting samples are not memorized during training, which is likely to happen in very few training epochs [10, 3], especially considering that the very few training bias-conflicting samples are more likely to be overfitted, as reported in [8, 25]. One could argue that an annotated validation set could be used for properly regularizing the bias amplification model, but the latter cannot be assumed available a priori in real cases, this being en emerging argument of discussion among the community [10, 3].
For this reason, in this work and differently from the previously cited research, we propose that syntethic bias-aligned samples could be used instead of the original training set, solving the bias-conflicting memorization issue by construction. However, an unsupervised real-world scenario poses a problem for generating a pure distribution of bias-aligned samples. As a solution, we propose to leverage diffusion models for generating a synthetic bias-aligned training set to be used for training a bias amplifier, inspired by recent works, showing how they can manifest distorted, unfair, or stereotypical representations deviating from the true underlying data distribution [26], when exposed to biases present in the training data  [15]. While the vast majority of available works regarding generative models and debiasing refer to strategies for obtaining unbiased synthetic data generation [27, 28], up to our knowledge, we are the first to propose the usage of generative models for bias amplification. The closest available work, employing generative models for model debiasing, is Jung et al. [29], where the authors introduce a debiasing method exploiting a GAN [30] trained for style-transfer between several biased appearances, with a different objective. Here, a target model is debiased through contrastive learning between images affected by different biases for bias-invariant representations.
@@
In our work, we show that it is possible to leverage the intrinsic tendency of diffusion models to capture bias in data, using conditioned generations to infer the biased distribution of each target class, thus allowing the sampling of new synthetic images, sharing the same bias pattern of original train data. Our method provides a memorization-free bias amplification model, capable of providing precise supervisory signals for accurate model debiasing virtually compatible with any debiasing approach involving auxiliary models.

### v4

@@ Abstract
The effectiveness of deep learning models in classification tasks is often challenged by the quality and quantity of training data whenever they are affected by strong spurious correlations between specific attributes and target labels. This results in a form of bias affecting training data, which typically leads to unrecoverable weak generalization in prediction. This paper addresses this problem by leveraging bias amplification with generated synthetic data only: we introduce Diffusing DeBias (DDB), a novel approach acting as a plug-in for common methods of unsupervised model debiasing, exploiting the inherent bias-learning tendency of diffusion models in data generation. Specifically, our approach adopts conditional diffusion models to generate synthetic bias-aligned images, which fully replace the original training set for learning an effective bias amplifier model to be subsequently incorporated into an end-to-end and a two-step unsupervised debiasing approach. By tackling the fundamental issue of bias-conflicting training samples’ memorization in learning auxiliary models, typical of this type of technique, our proposed method outperforms the current state-of-the-art in multiple benchmark datasets, demonstrating its potential as a versatile and effective tool for tackling bias in deep learning models.
@@ 1 Introduction
Deep learning models have shown impressive results in various vision tasks, including image classification. However, their success is highly dependent on the quality and representativeness of the training data. When a deep neural network is trained on a biased dataset, so-called “shortcuts”, corresponding to spurious correlations (i.e., biases) between irrelevant attributes and labels, are learned instead of semantic class attributes, affecting model generalization and impacting test performances. In other words, the model becomes biased towards specific training sub-populations presenting biases, known as bias-aligned, while samples not affected by the bias are referred to as bias-conflicting or unbiased. In this scenario, where it is generally assumed that most of the training samples available in a biased data set are biased (typically, $95\%$ or higher nam2020learning ), several debiasing methods have been proposed addressing the problem from different points of view NEURIPS2022_75004615_LWBC ; tartaglione2023disentangling ; pastore2024lookingmodeldebiasinglens ; kim2021biaswap ; Lee_Park_Kim_Lee_Choi_Choo_2023 . When annotations on bias attributes are available, common supervised debiasing strategies involve reweighting the training set to give more importance to the few available bias-conflicting data Lee_Park_Kim_Lee_Choi_Choo_2023 , for example, by upsampling li2019repair or upweighting Sagawa*2020Distributionally them. Conversely, unsupervised debiasing methods, which assume no prior availability of bias information, typically involve an auxiliary model to estimate and convey information on bias for reweighting the training set, employing its signal to guide the training of a debiased target model, while diminishing its tendency to learn bias shortcuts.
@@
To this end, there are several existing works nam2020learning ; NEURIPS2022_75004615_LWBC ; Lee_Park_Kim_Lee_Choi_Choo_2023 ; sohoni2020no exploiting intentionally biased auxiliary models, trained to overfit bias-aligned samples, thus being very confident in both correctly predicting bias-aligned training samples as well as misclassifying bias-conflicting ones Lee_Park_Kim_Lee_Choi_Choo_2023 . Such a distinct behavior between the two subpopulations is required to convey effective bias supervisory signals for target model debiasing, including gradients or per-sample loss values, which, e.g., should ideally be as high as possible for bias-conflicting and as low as possible for bias-aligned nam2020learning . However, such desired behavior is severely hindered by bias-conflicting training samples, despite their low number. Indeed, the low cardinality of this population impedes a suitable modeling of the correct data distribution, yet, they are sufficient to harm the estimation of a reliable bias-aligned data distribution, despite the latter having much higher cardinality. This scenario occurs under two main conditions. First, bias-conflicting samples in the training set act as noisy labels, affecting the bias learning for the auxiliary models Lee_Park_Kim_Lee_Choi_Choo_2023 . Second, these models tend to overfit and memorize bias-conflicting training samples within very few epochs, failing to provide different learning signals for the two subpopulations zarlenga2024efficient . Despite the available strategies, including the usage of the Generalized Cross Entropy (GCE) loss function nam2020learning ; pastore2024lookingmodeldebiasinglens , ensembles of auxiliary models NEURIPS2022_75004615_LWBC ; Lee_Park_Kim_Lee_Choi_Choo_2023 , or even ad hoc solutions tailored to the specific dataset considered (e.g., training for only one epoch or assuming the availability of a bias-annotated validation pmlr-v139-liu21f_JTT ), how to define a protocol for effective and robust auxiliary model training for unsupervised debiasing remains an open issue.
@@
Remarkably, recent works pastore2024lookingmodeldebiasinglens ; Lee_Park_Kim_Lee_Choi_Choo_2023 ; NEURIPS2022_75004615_LWBC underline how the target model generalization is profoundly impacted by the protocol used to train such an auxiliary model. In principle, if one could remove all bias-conflicting samples from the training set, it would be possible to train a bias-capturing model robust to bias-conflicting samples’ memorization and interference, and capable of conveying ideal learning signals Lee_Park_Kim_Lee_Choi_Choo_2023 .
@@
To this aim, our intuition is to elude these drawbacks by skipping the use of the actual training data in favor of another set of samples synthetically generated by an adequately trained diffusion model. Specifically, we propose to circumvent bias-conflicting interference on auxiliary models leveraging synthetic biased data obtained by exploiting Conditional Diffusion Probabilistic Models (CDPMs). We aim to learn a per-class biased image distribution, from which we sample an amplified synthetic bias-aligned subpopulation. The resulting synthetic bias-aligned samples are then exploited for training a Bias Amplifier model, instead of using the original actual training set, solving the issue of bias-conflicting overfitting and interference by construction, as the bias-capturing model does not see the original training set at all. We refer to the resulting approach as Diffusing DeBias (DDB). The intuition is indirectly supported by recent works studying the problem of biased datasets for image generation tasks d2024openbias ; kim2024training , and showing how diffusion models’ generations can be biased, as a consequence of bias present in the training sets (See Sec. 2).
@@
This tendency can be seen as an undesirable behavior of generative models, which affects the fairness of generated images, and recent articles have started to propose solutions to decrease the inclination of diffusion models to learn biases DBLP:journals/corr/abs-2408-15094 ; kim2024training ; DBLP:conf/icb/PereraP23 . However, we claim that what is considered a drawback for generative tasks can be turned into an advantage for image classification’s unsupervised model debiasing, as it allows the training of a robust auxiliary model that can ideally be plugged into various debiasing schemes, solving bias-conflicting training samples memorization by construction. For this reason, we refer to DDB as a plug-in for debiasing methods in image classification. To prove its effectiveness and versatility, we incorporate DDB’s bias amplifier into two different debiasing frameworks, which we refer to as Recipe I and Recipe II, including a two-step and an end-to-end approach (see Sec. 2). Specifically, we exploit our bias-amplifier to: (i) extract subpopulations further used as pseudo-labels within the popular G-DRO Sagawa*2020Distributionally algorithm (Recipe I, Sec. 3.3.1), and (ii) provide a loss function for the training set to be used within an end-to-end method (Recipe II, Sec. 3.3.2). Both approaches are competitive with the state-of-the-art on popular benchmark datasets. In summary, our main contributions can be regarded as follows:
@@
  • We introduce DDB, a novel and effective debiasing framework that exploits CDPMs to learn per-class bias-aligned distributions in biased datasets with unknown bias information. The resulting CDPM is used to generate synthetic bias-aligned images later used to train a robust bias amplifier model, thus avoiding real bias-conflicting training sample memorization and construction interference (Sec. 3).
@@
  • We propose the usage of DDB as a versatile plug-in for debiasing approaches, designing two recipes: one based on a two-step procedure (Sec. 3.3.1), and the other as an end-to-end (Sec. 3.3.2) algorithm. Both proposed approaches prove to be effective, achieving state-of-the-art results (Sec. 4) across popular benchmark datasets, characterized by different types of single or multiple biases.
@@ 2 Related Work
In this section, we provide a broad categorization of model debiasing works.
@@
Supervised Debiasing. Supervised methods exploit available prior knowledge about the bias (e.g., bias attribute annotations), indicating whether a sample exhibits a particular bias (or not). A common strategy consists of training a bias classifier to predict such attributes so that it can guide the target model to extract unbiased representations alvi2018turning ; Xie2017ControllableIT . Alternatively, bias labels can be thought of as pre-defined encoded subgroups among target classes, allowing for the exploitation of robust training techniques such as G-DRO Sagawa*2020Distributionally . At the same time, bias information can be employed to apply regularization schemes aiming at forcing invariance over bias features barbano2023unbiased ; tartaglione2021end .
@@
Unsupervised Debiasing. It considers bias information unavailable (the most likely scenario in real-world applications), thus having the highest potential impact among the described paradigms. Among such methods, a popular strategy consists in involving auxiliary models at any level of the debiasing process. We categorize these methods into two-step  pmlr-v139-liu21f_JTT ; sohoni2020no ; kim2021biaswap , and end-to-end  nam2020learning ; li2022discover , in the following paragraphs.
@@
End-to-end approaches. Here, bias mitigation is performed online with dedicated optimization objectives. Employing techniques such as ensembling and noise-robust losses has been proposed, though they require careful tuning of all the hyperparameters, which may require the availability of validation sets with bias annotations to work. In Learning from Failure (LfF), Nam et al. nam2020learning assume that bias features are learned earlier than task-related ones: they employ a bias-capturing model to focus on easier (bias-aligned) samples through the GCE loss zhang2018gce . At the same time, a debiased network is jointly trained to assign larger weights to samples that the bias-capturing model struggles to discriminate. In DebiAN, Li et al. li2022discover jointly train an auxiliary model (the discoverer) and a classifier (the target model) in an alternate scheme so that the discoverer can identify bias learned by the classifier during training, optimizing an ad-hoc loss function for bias mitigation.
@@
Two-step methods. Here, the auxiliary model is employed in a first step to identify bias in training data, generally through pseudo-labeling, while the second stage consists of actual bias mitigation exploiting the inferred pseudo-labels. In pmlr-v139-liu21f_JTT , the auxiliary model is trained with standard ERM and then used in inference on the training set, considering misclassified samples as bias-conflicting and vice-versa: debiasing is brought on by up-sampling the predicted bias-conflicting samples. Notably, they must rely on a validation set with bias annotations to stop the auxiliary model’s training after a few epochs to avoid memorization and to tune the most effective upsampling factor. In NEURIPS2022_75004615_LWBC , a set of auxiliary classifiers is ensembled (bootstrap ensembling) into a bias-commmittee, proposing that debiasing can be performed using a weighted ERM, where the weights are proportional to the number of models in the ensemble misclassifying a certain sample. Finally, a recent trend involves the usage of vision-language models to identify bias, further exploiting bias predictions to mitigate biases in classification models eyuboglu2022domino ; kim2024discovering ; zhang2023diagnosing .
@@
Among the described methods, some exploit the auxiliary model for bias amplification (Lee_Park_Kim_Lee_Choi_Choo_2023 ; nam2020learning ; sohoni2020no ) under the general assumption that a bias amplification model would either misclassify or be less confident on bias-conflicting samples. In this case, the final debiasing performance is heavily dependent on such behavior, and specifically, on how strongly the auxiliary model captures the bias in training data pastore2024lookingmodeldebiasinglens ; Lee_Park_Kim_Lee_Choi_Choo_2023 . However, this fundamental assumption only holds if bias-conflicting samples are not memorized during training, which is likely to happen in very few training epochs zarlenga2024efficient ; pastore2024lookingmodeldebiasinglens , especially considering that the very few training bias-conflicting samples are more likely to be overfitted, as reported in Sagawa*2020Distributionally ; cao2019learning . One could argue that an annotated validation set could be used for properly regularizing the bias amplification model, but the latter cannot be assumed available a priori in real cases, being this an emerging argument of discussion among the community zarlenga2024efficient ; pastore2024lookingmodeldebiasinglens . For this reason, differently from the previously cited research, we propose in this work that synthetic bias-aligned samples could be used instead of the original training set, solving the bias-conflicting memorization issue by construction. However, a real-world scenario using data with unknown bias poses a problem for generating a pure distribution of bias-aligned samples. As a solution, we propose to leverage diffusion models for generating a synthetic bias-aligned training set to be used for training a bias amplifier.
While the vast majority of available works regarding generative models and debiasing refer to strategies for obtaining unbiased synthetic data generation bhat2023debiasinggenerativemodelsusing ; gerych2023debiasing , to the best of our knowledge, we are the first to propose the use of generative models for obtaining an amplified biased distribution.
The closest work employing generative models for model debiasing is Jung et al. jung2023fighting , where a debiasing method exploiting a GAN 10.1145/3422622 is introduced, which is trained for style-transfer between several biased appearances with different objectives. Here, a target model is debiased through contrastive learning between images affected by different biases for bias-invariant representations. Similar strategies, where bias-conflicting samples are augmented to support model debiasing, can be found in a few other works biasadv ; hwang2022selecmix ; dataaug1 . In dataaug1 , the problem of bias-conflicting memorization is stressed, proposing a bias-unsupervised strategy to augment and increase the diversity of bias-conflicting features, exploiting a bias amplifier analogously to nam2020learning for their identification, while improving the generalization of a target debiased model. In biasadv , an intentionally biased auxiliary model is exploited, which is trained on the biased dataset, supplementing bias-conflicting samples for the target debiased model. Here, images are generated by solving an optimization problem consisting of attacking the auxiliary model predictions while preserving the ones of the target debiased model. In hwang2022selecmix , bias-conflicting samples are again augmented exploiting the popular mix-up approach zhang2018mixupempiricalriskminimization . Specifically, an auxiliary model is employed to amplify easy-to-learn features, and mix-up pairs with either the same ground truth label, but different biased features (evaluated with clustering in the feature space of an intentionally biased model), or samples with different biased features (i.e., belonging to different clusters), but showing the same ground truth label.
@@
Differently from the previously described approaches, we are proposing an alternative strategy to avoid bias-conflicting memorization, while significantly improving the generalization of the target debiased model. In our work, we show how it is possible to leverage the intrinsic tendency of diffusion models to capture bias in data, using conditioned generations to infer the biased distribution of each target class, thus allowing the sampling of new synthetic images, sharing the same bias pattern of the original training data. Our method designs a memorization-free bias amplified model, capable of providing precise supervisory signals for accurate model debiasing, virtually compatible with any debiasing approach involving auxiliary models.